OpenStudy (anonymous):

Please help Translate the following verbal statement into an algebraic equation and the solve: Use x for your variable. The quotient of the sum of a number and nine and seven times the number is nine.

OpenStudy (jhulian):

(x+9)/7x = 9 x + 9 = 63x x = 9/62